a medivac helicopter travels 78 km due south of its base to pick up a patient the helicopter then travels 93 km due north to a hospital the entire trips take 1.22h what is the avarage speed and the average velocity.
Answer by nerdybill(7012)

a medivac helicopter travels 78 km due south of its base to pick up a patient the helicopter then travels 93 km due north to a hospital the entire trips take 1.22h what is the avarage speed and the average velocity.
.
total distance traveled = 78+93 = 171 km
.
average speed = 171/1.22 = 140.2 km/hr